OpenStudy (anonymous):

Find the rational roots of x^4+8x^3+7x^2-40x-60=0

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

You can use the rational roots theorem to guess and check the possible roots. You will find that (x + 2) and (x + 6) are factors of this polynomial. The other factor is (x^2 - 5). Can you obtain the rational roots from here?

OpenStudy (anonymous):

yes but how did you get those numbers

OpenStudy (anonymous):

Using the Rational Roots Theorem ( http://www.mathwords.com/r/rational_root_theorem.htm) you guess and check the possible values. The possible rational roots are +- p/q where p are the factors of 60 and q are the factors of 1. I.e., the possible rational roots are +- 1, +-2, +-4, +-5, +-6., etc... By guess and check we can try x = 1, x = -1, x = -2, x = 2 and x = -2. We find that x = -2 results in a 0. That means (x + 2) is a factor. Divide the original polynomial by (x + 2) to obtain a simpler polynomial of degree 3 and repeat the same steps to obtain the next factor.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I repeated x = -2 twice by accident in the explanation above. Anyway, to show x = -2 results in a 0: (-2)^4+8(-2)^3+7(-2)^2-40(-2)-60 = 16 - 64 + 28 + 80 - 60 = 0 So if x = -2, that means (x + 2) must be a factor of the original polynomial. Division yields (see attached file). Our new polynomial is x^3 + 6x^2 - 5x - 30. We can use the same Rational Roots Theorem to determine the remaining possible rational root(s).